# Aspects of the bulk flat space limit in AdS/CFT

### v1

@@ VII Discussion
When following the same strategy for massless particles, we learn that to generate a finite energy particle, one must consider a double scaling limit (56) in which the AdS scale $\Lambda\to\infty$, the boost rapidity $\eta\to\infty$ keeping $\Lambda^{-1}\,e^{\eta}$ fixed. The resulting wave functions do not lead to plane waves, they involve several energies, depend on arbitrary null vectors and remember the AdS conformal dimension $\Delta$ in their profile. These wave functions resemble similar quantities considered in euclidean AdS and look similar to some of the conformal primary wave functions considered in the celestial holography literature [31, 32]. Our preliminary investigation on these wave functions suggests one may be able to extract a massless S-matrix using a Schwinger parameterization that is reminiscent of the types of Mellin transforms that appeared in celestial holography. Our wave functions still allow an easy operational AdS state preparation in terms of Euclidean operator boundary insertions. In the massless limit, our state preparation recovers the naïve geometric picture that emerged in a HKLL approach involving real time physics in a narrow band around $t=\pm\frac{\pi}{2}$, but the full details are different.
There are multiple directions for future research. First, we would like to understand the massless case in detail. Our massless waveform (57) and their Schwinger parameterization (58) must be related to the basis of massless scalar conformal primaries discussed in celestial holography (see [34, 41, 42], for example)
@@
A fourth point to be made is that this work explored the flat space limit in the bulk from a representation theory perspective. If we were to take the same perspective on the dual boundary theory, the literature describes the corresponding Ìnönü-Wigner contraction as a Carrollian limit (see [43] for a review). Relating the bulk wave functions derived in this work with their boundary counterparts may provide a constructive perspective on the meaning of the Carrollian limit and the physics of Carrollian quantum field theories.
Finally, our entire discussion was concerned with vacuum physics of a few particles in the regime where QFT on a fixed curved spacetime makes sense. What about thermal states ? Clearly, the flat Schwarzschild black hole is the flat limit of the small AdS black hole. Just as the latter are thermodynamically unstable, black holes in flat space decay too. It would be interesting to explore whether Carrollian physics is responsible for such instabilities, extending results in 3d pure gravity [44].

### v2

@@ VII Discussion
When following the same strategy for massless particles, we learn that to generate a finite energy particle, one must consider a double scaling limit (56) in which the AdS scale $\Lambda\to\infty$, the boost rapidity $\eta\to\infty$ keeping $\Lambda^{-1}\,e^{\eta}$ fixed. The resulting wave functions do not lead to plane waves, they involve several energies, depend on arbitrary null vectors and remember the AdS conformal dimension $\Delta$ in their profile. These wave functions resemble similar quantities considered in euclidean AdS and look similar to some of the conformal primary wave functions considered in the celestial holography literature [Raclariu:2021zjz, Pasterski:2021rjz]. Our preliminary investigation on these wave functions suggests one may be able to extract a massless S-matrix using a Schwinger parameterization that is reminiscent of the types of Mellin transforms that appeared in celestial holography. Our wave functions still allow an easy operational AdS state preparation in terms of Euclidean operator boundary insertions. In the massless limit, our state preparation recovers the naïve geometric picture that emerged in a HKLL approach involving real time physics in a narrow band around $t=\pm\frac{\pi}{2}$, but the full details are different.
There are multiple directions for future research. First, we would like to understand the massless case in detail. Our massless waveform (57) and their Schwinger parameterization (58) must be related to the basis of massless scalar conformal primaries discussed in celestial holography (see [Pasterski:2016qvg, Pasterski:2017kqt, Cotler:2023qwh], for example)
@@
A fourth point to be made is that this work explored the flat space limit in the bulk from a representation theory perspective. If we were to take the same perspective on the dual boundary theory, the literature describes the corresponding Ìnönü-Wigner contraction as a Carrollian limit (see [Bagchi:2025vri] for a review). Relating the bulk wave functions derived in this work with their boundary counterparts may provide a constructive perspective on the meaning of the Carrollian limit and the physics of Carrollian quantum field theories.
Finally, our entire discussion was concerned with vacuum physics of a few particles in the regime where QFT on a fixed curved spacetime makes sense. What about thermal states? Clearly, the flat Schwarzschild black hole is the flat limit of the small AdS black hole. Just as the latter are thermodynamically unstable, black holes in flat space decay too. It would be interesting to explore whether Carrollian physics is responsible for such instabilities, extending results in 3d pure gravity [Aggarwal:2025hji].